A car is traveling with a velocity of 14 m/s when the driver accelerates it at a rate of 2.3m/s 2 for 8
seconds. can you explain me step by step how to do it.
1
Expert's answer
2020-11-06T10:14:38-0500

Let's find the velocity and distance travelled by the car in "t = 8s". By definition, the acceleration is:


"a = \dfrac{v_f - v_i}{t}"

where "v_i = 14m\/s" is car's initial speed and "v_f" is car's final speed after time "t". Expressing the final speed, obtain:


"v_f = v_i + at = 14m/s + 2.3m/s^2\times 8s = 32.4m/s"

The distance travelled by the car is:


"d = v_it + \dfrac{at^2}{2} = 14m/s\times 8s + \dfrac{2.3m/s^2\times (8s)^2}{2} = 112m + 73.6m = 185.6m"

Answer. Distance 185.6m, speed 32.4m/s.